OpenStudy (sophadof):

Find the point(s) of intersection of the line X + Y = 5 and the circle x2 + y2 = 25 by solving the system of equations.

OpenStudy (anonymous):

My geometry is very rusty, but I'll give it a shot... If I remember correctly, you will need to put both equations into slope-intercept forum. Doing this, you will get y= -x+5 and y= x+5 ( someone may want to double check this). After this, you use trace the two equations and will get an intersection of y=5. Again, my geometry is rusty and someone may want to double check.

OpenStudy (jtvatsim):

Iehmad is close. The idea is to solve the first equation for y. 1) This gives y = -x + 5. 2) You will then plug this into the second equation x^2 + y^2 = 25, to get x^2 + (-x + 5)^2 = 25. 3) Solve this for x. 4) You can then use the x's you find to solve for y. 5) Your final answer will be coordinates (x,y).

OpenStudy (sophadof):

wait..so @jtvatsim after i plug it in the step after would be x+(5)^2=25?

OpenStudy (jtvatsim):

The second equation is x^2 + y^2 = 25. We will replace y with y = -x + 5. That makes the equation: x^2 + (-x + 5)^2 = 25. Be careful to factor the entire parentheses. You should get x^2 + (-x + 5)*(-x + 5) = 25 which simplifies to x^2 + (x^2 - 5x - 5x + 25) = 25, x^2 + x^2 - 10x + 25 = 25. Does that make sense?

OpenStudy (jtvatsim):

Of course, there is still work to do from where I left off, but hopefully that gets you started. :)

OpenStudy (jtvatsim):

You can check your answer against this graph I did on Desmos.com

OpenStudy (sophadof):

I am completely lost :( @jtvatsim

OpenStudy (sophadof):

it makes sense..it's just moving on from what you did is losing me.. @jtvatsim

OpenStudy (jtvatsim):

No problem. Here's how I would do it. After simplifying the work I did above you should get: 2x^2 - 10x + 25 = 25. Did you get that?

OpenStudy (sophadof):

yes!!(: @jtvatsim i actualy did haha

OpenStudy (jtvatsim):

OK, great! Now, let's subtract 25 from both sides leaving us with: 2x^2 - 10x = 0

OpenStudy (jtvatsim):

I can factor out a 2 (not essential, but why not?) to get 2(x^2 - 5x) = 0 then, x^2 - 5x = 0 (after dividing by 2). Factoring, gives me x(x - 5) = 0. so x = 0, or x = 5.

OpenStudy (jtvatsim):

Now WAY BACK in the beginning, we said that y = -x + 5. So we can plug in our new found values of x to get the y values. x = 0, means y = -0 + 5 = 5. This is the point (0,5). x = 5, means y = -5 + 5 = 0. This is the point (5,0). We are done. Yay! :)

OpenStudy (jtvatsim):

Any questions? :)

OpenStudy (sophadof):

haha hang on..just gonna try it(: @jtvatsim

OpenStudy (jtvatsim):

Sounds good.

OpenStudy (sophadof):

Thank you for staying in this with me though!!!!(: @jtvatsim

OpenStudy (jtvatsim):

No worries, glad to help if I can. :)

OpenStudy (sophadof):

yeah i got x=5 @jtvatsim

OpenStudy (jtvatsim):

And you should also get x = 0 as another solution.

OpenStudy (sophadof):

oh! @jtvatsim gotcha!

OpenStudy (jtvatsim):

Great! If you had done something like x^2 = 5x and divided by x, you would get x = 5, but dividing by x accidentally causes you to miss the x = 0 solution. Just wanted to be sure you saw that. :)

OpenStudy (sophadof):

yes thank you!!!!! @jtvatsim I hate to ask you this but...do you have email, I have another math question and its with triangles..

OpenStudy (jtvatsim):

No worries, feel free to post it on OpenStudy just in case I have to run here soon. At least someone else would be able to jump in. :)

OpenStudy (sophadof):

yes thank you!!!!! @jtvatsim I hate to ask you this but...do you have email, I have another math question and its with triangles..

OpenStudy (sophadof):

@jtvatsim

OpenStudy (jtvatsim):

And we are to find the value of x correct?

OpenStudy (sophadof):

yes

OpenStudy (jtvatsim):

Alright, first question, do you know what the ~ symbol means?

OpenStudy (sophadof):

similar right?

OpenStudy (jtvatsim):

Correct!

OpenStudy (jtvatsim):

So, that basically means the triangles have the same shape (maybe not the same size).

OpenStudy (jtvatsim):

The problem is that the picture didn't really do a good job at showing this (especially since one is upside down!). I'll go ahead and fix that.

OpenStudy (sophadof):

haha thank you

OpenStudy (jtvatsim):

 That's better.

OpenStudy (jtvatsim):

Now, whenever we have similar triangles, we like to notice the sides that match and set up a "proportion". Something like these fractions: (big triangle side 1)/(big triangle side 2) = (small triangle side 1)/(small triangle side 2) Make sense so far?

OpenStudy (sophadof):

yes!

OpenStudy (jtvatsim):

Cool! So, now let's pick some sides! What would you like to pick for "big triangle side 1"? You can pick whatever you like (based on the drawing of course). :)

OpenStudy (sophadof):

ill choose 75

OpenStudy (jtvatsim):

OK, so right now we have: 75/(big triangle side 2) = (small triangle side 1)/(small triangle side 2). We need to keep going though. I would suggest that we pick "small triangle side 1" now. We need to pick the same part of the smaller triangle. (In this case the bottom). What would that be for the small triangle?

OpenStudy (sophadof):

4x-10

OpenStudy (jtvatsim):

Perfect!

OpenStudy (jtvatsim):

We now just repeat the process for the other sides we know. We should get something like this...

OpenStudy (jtvatsim):

75/30 = (4x - 10)/(x + 5) Looks a little scary, but we have now got the question ready to solve. Any questions so far before we go on?

OpenStudy (sophadof):

nope im good!!

OpenStudy (jtvatsim):

Very good! :)

OpenStudy (jtvatsim):

Right now, cross multiplication is your friend (or if you prefer multiply each fraction by the bottoms) and this gives us: 75*(x + 5) = 30*(4x - 10)

OpenStudy (jtvatsim):

Very important step right there, are you ok with that?

OpenStudy (sophadof):

yes! then do i solve?

OpenStudy (jtvatsim):

Yes! Now you have a "nice" looking equation (you do have some distributing to do, so be careful).

OpenStudy (sophadof):

so...75x+375=120x-300?

OpenStudy (jtvatsim):

Yes, that is correct!

OpenStudy (sophadof):

yay!!!!!!!!!(: thank you!

OpenStudy (jtvatsim):

no problem!

OpenStudy (sophadof):

then subtract 75 from 120?

OpenStudy (jtvatsim):

Sure, that leaves you with 45x I believe.

OpenStudy (sophadof):

x=15?(:

OpenStudy (jtvatsim):

You got it!

OpenStudy (sophadof):

yay!!!!! Thank you soooo much!!(: do i ned to plug in x for those equations or no?

OpenStudy (jtvatsim):

Only if the question wanted you to find the actual sides of the unknown triangle. Sometimes they do want you to find that.

OpenStudy (sophadof):

oh alrighty!!! thank you so much!!

OpenStudy (jtvatsim):

That's easy to find though based on the information we now have.

OpenStudy (jtvatsim):

Your welcome!

OpenStudy (sophadof):

Have a good night? day? haha

OpenStudy (jtvatsim):

More like evening. :) You too!